# Neuromorphic Quantum Neural Networks with Tunnel-Diode Activation Functions

### v1

@@ III Results
The neural network architecture shown in Fig. 1 was trained through 10 epochs with a learning rate of 0.01 and a batch size of 256. The MNIST training data was randomly split into 80% training and 20% validation, and the accuracy and loss values after each epoch were recorded for the training and validation sets. This training procedure was applied to the three topologically identical networks that utilised ReLU, sigmoid and TDAF, respectively. To account for variations caused by the random initialisation, the entire experiment was repeated 20 times. For each activation function, the means of the losses and accuracies after each epoch across the 20 repeats was recorded. All computations were performed on an Apple Macbook Pro M3 using an 8-core CPU.
@@ III-B Accuracy on Test Dataset
Each model was trained 20 times to reduce statistical noise caused by the initialisation and to measure the deviations between different runs. Figure 5 demonstrates the performance of the models on the test set after training, showing the median and interquartile range. TDAF achieves a mean accuracy of $96.14\%$ on the test set, compared to $95.32\%$ for ReLU and $94.44\%$ for sigmoid. This result is similar to the final accuracies of the validation set of $95.86\%$, $95.04\%$ and $94.13\%$ for TDAF, ReLU and sigmoid, respectively. Thus, outperforming the other activation functions based on the median and mean test accuracy, TDAF also exhibits a smaller interquartile range and overall range indicating more stable training dynamics. Additionally, TDAF outperforms both Sigmoid and ReLU in terms of the loss and accuracy throughout training as well as the performance on the test set after training.
@@
Since TDAF has demonstrated superior characteristics compared to other activation functions, we assess the performance of the TDAF-based neural network across the ten classes of the MNIST dataset. Figure 6 presents the loss during training, the confusion matrix and the accuracy across different classes. The class accuracies range from $92.73\%$ for the digit 2 to $98.07\%$ for the digit 4. The most frequent misclassification was for the digits 2 being identified as a digit 8, occurring 42 times—an outcome that aligns with empirical observations [50]. It is noteworthy that the strong performance observed in this study, despite the simplicity of the FNN model, underscores the advantages conferred by TDAF, demonstrating its potential to enhance classification accuracy in neural networks with minimal architectural complexity.
IV Discussion
@@
The successful adaption of the characteristic I-V curve of a tunnel diode as an activation function for neural networks raises interesting questions on the underlying mechanisms causing the improved performance and potential future applications. TDAF exhibited superior training dynamics, stability and accuracy on the MNIST dataset compared with the sigmoid and ReLU activation functions. This suggests that the TDAF creates a simpler loss landscape when compared to ReLU and sigmoid.
@@
In practice, a TDAF-based neural network can be implemented using readily available commercial tunnelling [23] and resonant tunnelling [24] diodes. It is noteworthy that systems based on tunnelling diodes consume significantly less power than traditional integrated electronic circuits [26], presenting an opportunity for developing semiconductor chips optimised for power-constrained applications, such as autonomous vehicles, where onboard energy resources are highly limited [9]. Furthermore, NC schemes that exploit negative differential resistance—a key characteristic of tunnelling diodes—have already been proposed [25, 22], suggesting that TDAF functionality could be integrated into existing systems.
Alternatively, a TDAF can be implemented using scanning tunnelling microscopy (STM) instrumentation [28]. STM amplifiers operate at microwave frequencies, providing high signal-to-noise ratios and enabling differential resistance spectroscopy measurements [51], which can be used to compute the derivative of TDAF. In low-frequency designs, this derivative can also be obtained using simple analog electronic circuits [52]. Furthermore, QT of individual electrons has been experimentally observed in quantum dots [25], which serve as fundamental building blocks for quantum NC systems and can therefore be integrated with them [10].
AI will be pivotal in future space exploration, but significant technological challenges must be addressed, particularly in energy consumption and thermal management [53]. Modern AI systems rely on efficient cooling to dissipate heat from graphics processing units, yet the vacuum of space renders conventional air-based cooling methods ineffective [53]. This makes heat dissipation a critical concern for spacecraft, satellites and space exploration vehicles, necessitating innovative cooling solutions to ensure the reliable operation of sensitive electronics in extreme environments [54].
@@
V Conclusion
We introduced a novel activation function, the tunnel-diode activation function (TDAF), derived from the current–voltage (I–V) characteristic of a tunnel diode. TDAF demonstrated superior performance compared to two widely used activation functions—ReLU and sigmoid—achieving higher accuracy and lower loss during both training and evaluation. Furthermore, TDAF exhibited stable behaviour throughout training, highlighting its robustness across different machine learning tasks.
We suggested that the improved performance of TDAF can be attributed to the QT effect that underpins the operation of tunnel diodes. This effect introduces greater nonlinearity and expands data into a broader set of nonlinear functionals, enabling ML algorithms to process information more efficiently. This physically grounded approach opens new avenues for integrating quantum-inspired mechanisms into ML models, bridging the gap between AI and emerging neuromorphic computing paradigms.
Future research will focus on further theoretical exploration of the QT-based nonlinearity of TDAF and its implications for network dynamics. Additionally, practical implementation of TDAF in hardware-based neuromorphic circuits will be pursued, potentially enabling the development of energy-efficient, high-speed AI systems capable of operating in environments unsuitable for conventional or qubit-based quantum computing hardware.
Comparing the performance of TDAF implementations across different neural network architectures—such as recurrent and convolutional neural networks, RC systems and transformer-based models—will further test the robustness of the activation function and establish which types of tasks TDAF performs best. Moreover, studying how activation functions affect the loss landscape will provide insight into whether other physically inspired functions may offer advantages for ML applications.
As previously mentioned, our results indicate that implementing tunnel diode circuits, which take input data encoded as currents and process it analogously to a feedforward neural network, may yield promising outcomes. Pursuing this research direction could lead to cost-effective AI systems with improved computational efficiency, bridging the gap between quantum-inspired NC systems and practical ML implementations.

### v2

@@ III Results
The neural network architecture shown in Fig. 1 was trained through 10 epochs with a learning rate of 0.01 and a batch size of 256. The MNIST training sub-dataset was randomly split into 80% training and 20% validation, and the accuracy and loss values after each epoch were recorded for the training and validation sets. This training procedure was applied to the three topologically identical networks that utilised ReLU, sigmoid and TDAF, respectively. To account for variations caused by the random initialisation, the entire experiment was repeated 20 times. For each activation function, the means of the losses and accuracies after each epoch across the 20 repeats was recorded. All computations were performed on an Apple Macbook Pro M3 using an 8-core CPU.
@@ III-B Accuracy on Test Dataset
Figure 4 demonstrates the performance of the models on the test set after training, showing the median and interquartile range. TDAF achieves a mean accuracy of $96.14\%$ on the test set, compared to $95.32\%$ for ReLU and $94.44\%$ for sigmoid. This result is similar to the final accuracies of the validation set of $95.86\%$, $95.04\%$ and $94.13\%$ for TDAF, ReLU and sigmoid, respectively. Thus, outperforming the other activation functions based on the median and mean test accuracy, TDAF also exhibits a smaller interquartile range and overall range indicating more stable training dynamics. Additionally, TDAF outperforms both sigmoid and ReLU in terms of the loss and accuracy throughout training as well as the performance on the test set after training.
@@
Since TDAF has demonstrated superior characteristics compared to other activation functions, we assess the performance of the TDAF-based neural network across the ten classes of the MNIST dataset. Figure 5 presents the loss during training, the confusion matrix and the accuracy across different classes. The class accuracies range from $92.73\%$ for the digit 2 to $98.07\%$ for the digit 4. The most frequent misclassification was for the digits 2 being identified as a digit 8, occurring 42 times, an outcome that aligns with empirical observations [58]. It is noteworthy that the strong performance observed in this study, despite the simplicity of the feedforward neural network model, underscores the advantages conferred by TDAF, demonstrating its potential to enhance classification accuracy in neural networks with minimal architectural complexity.
IV Comparison with a Quantised Neural Network
While the class accuracies achieved in this work are already feasible, more advanced architectures, such as deep neural networks with extended hidden layers and convolutional neural networks, can reach even higher performance. In principle, the TDAF could be incorporated into these models; however, demonstrating state-of-the-art MNIST classification accuracy is not the objective of this article. Rather, our focus is on demonstrating that tunnel diodes, beyond their established applications in quantum-cognitive [28, 29, 30] and spiking [20, 17, 18] neural networks, offer a pathway towards inexpensive, low-power neuromorphic hardware [10]. This naturally invites comparison with neural-network frameworks designed explicitly for efficient, resource-constrained computation, with quantised neural networks being the most widely adopted approach [42, 43].
Quantisation (not to be confused with the quantum-mechanical effects that underpin the operation of tunnel diodes) in neural networks is the process of reducing the numerical precision of weights, activations or both from standard 32-bit floating-point values to lower-precision formats such as 8-bit integers, ternary values {1, $\{-1,\,0$,\,+1\} and binary values [42, 43]. The aim of quantisation is to make the network cheaper and faster to run on a system with limited computational power without significantly degrading accuracy.
Lower-precision numbers reduce memory footprint, cut data-movement costs (a major source of energy consumption) and enable the use of simpler, more efficient arithmetic hardware [59, 60]. This is essential for deploying deep learning models on edge devices, mobile processors, embedded systems and specialised accelerators where power, memory and computational resources are limited. Quantisation can also improve inference speed on general-purpose hardware, making it an important emerging machine learning technique [61].
To benchmark the performance of the TDAF-based neural network against a quantised model, we compare it with a customised ternary neural network [62]. Our implementation incorporates a balanced combination of different techniques drawn from contemporary ternary neural network research, including modern weight-quantisation strategies [63, 64, 65] and teacher–student training approaches [59].
In ternary neural networks, real-valued weights are mapped onto the discrete set {1, $\{-1,\,0$,\,+1\} using a ternarisation operator. Let $w$ denote a real-valued weight and $\Delta>0$ a threshold. The ternary weight $w^{t}$ is defined as
$w^{t}=\operatorname{Tern}(w)=\begin{cases}+1,&w>\Delta,\\[4.0pt] 0,&|w|\leq\Delta,\\[4.0pt] -1,&w<-\Delta\,.\end{cases}$ (11)
Certain ternary models employ scaled ternarisation [66] using a real-valued scaling factor $\alpha$, which captures the typical size of the significant weights, ignoring the small ones. This factor is introduced as
$w^{t}=\alpha\cdot\operatorname{Tern}(w),$ (12)
where
$\alpha=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}|w_{i}|,\qquad\mathcal{I}=\{\,i:|w_{i}|>\Delta\,\}\,.$ (13)
However, in this work we modify this step and ternarise the weights directly as $w^{t}=\operatorname{Tern}(w)$ (i.e. in our model, real-value scaling can yield adequate results; however, higher performance was achieved using fully ternarised weights).
We also use operator $\operatorname{Tern}$ as the activation function [63] as well as adopt a teacher–student training strategy, in which the forward activations and backpropagated gradients are computed using the ternarised weights $w^{t}$ and no bias term is included [59]. In this setup, however, the underlying full-precision weights remain part of the optimisation loop and are updated in full precision throughout training. Moreover, the output layer of the network relied on raw scores with an approach inspired by squared hinge loss and its derivative [63], enabling the final classification to be implemented directly via the argmax decision function. Another advantage of using the balanced ternary system is that, during inference, the activation function reduces to a simple sign operation [67].
However, because the ternarisation function in Eq. (11) is non-differentiable, gradient-based training requires an approximation to the derivative. This is typically achieved using the Straight-Through Estimator (STE) approach, which replaces the true derivative with an artificial mathematical expression during backpropagation [68]. A general STE is expressed as
$\frac{\partial w^{t}}{\partial w}\approx 1\,$, (14)
allowing the gradient to ‘pass through’ the quantisation step. We employ a more constrained version of the STE that limits gradient flow to a finite interval [69]
$\frac{\partial w^{t}}{\partial w}\approx\begin{cases}1,&|w|\leq 1,\\[4.0pt] 0,&|w|>1\,.\end{cases}$ (15)
In Table I, we compare the final MNIST classification accuracies of the TDAF, ReLU and sigmoid full-precision models discussed above, along with the ternary model. We emphasise that our goal was not to fully optimise any of the models for maximum performance reported in the literature (see, e.g., [66]); therefore, the reported accuracies should be interpreted accordingly. We also note that the input to the ternary model consisted of ternarised MNIST images.
We observe that the TDAF and ternary models achieve comparable accuracies. However, while the ternary model restricts the use of full-precision arithmetic, since such operations would be prohibitively expensive when implemented in hardware such as field-programmable gate arrays (FPGAs) [59], the TDAF model, as presented here, retains full-precision activations in the mathematical framework used as a testbench. However, in practical hardware implementations, where TDAF units would correspond to physical tunnel-diode elements, the activation function would be realised directly in inexpensive, low-energy analogue circuitry, eliminating the need for full-precision digital computation [10].
Figure 6: Convergence behaviour during training of the ternary neural network model used in this study.
It should be acknowledged that FPGAs are an established commercial technology but the use of analogue electronic devices in neuromorphic-style systems is still an emerging approach transitioning from laboratory prototypes to early industrial adoption. The TDAF model nevertheless can offer several potential advantages over the ternary model. First, implementing ternary representations on FPGAs requires additional workarounds, such as encoding trits using multi-bit binary patterns (e.g., 01, 10, etc. or similar schemes) to represent ternary values {1, $\{-1,\,0$,\,+1\}, which introduces overhead compared with native binary hardware [70]. Second, as also noted in [63], ternary models tend to converge more slowly than their full-precision counterparts (Fig. 6)—and therefore slower than the TDAF model, which in our experiments demonstrated faster convergence than both the ReLU- and sigmoid-based models. Third, as will be discussed in more detail below, inexpensive tunnel diodes can operate in harsh environments such as high altitude and space. However, commercial off-the-shelf FPGAs are not inherently radiation-hard and are vulnerable to single-event effects, which is especially critical for aerospace, high-altitude, nuclear and defence applications.
TABLE I: Model Accuracy Comparison (MNIST test sub-dataset)
Model Final accuracy
TDAF-DNN 95.86%
ReLU-DNN 95.04%
sigmoid-DNN 94.13%
ternary-DNN 94.35%
QT-ternary-DNNa 84.00%
  • a
    Hidden layers twice as small.
V TDAF: Origin of High Performance
V-A Analysis of Nonlinear Properties of TDAF
As discussed above, nonlinearity is essential for the operation of neural networks. In this section, we demonstrate that the increased nonlinearity of the TDAF plays a key role in its superior performance compared with traditional activation functions commonly used in machine learning. This discussion is warranted because earlier work on a tunnel-diode-based reservoir computing system [71] has also demonstrated that increasing the nonlinearity of the signal processing in a tunnel diode improves the overall performance of the neural network.
Analysing the degree of nonlinearity of any signal is a non-trivial task, as no standard metrics exist and nonlinearity itself arises from diverse physical and mathematical processes. For example, one method to compare the nonlinearity of two signals is to analyse their Taylor series expansions and compare the magnitudes of the higher-order terms. However, the Taylor series is an approximation and calculating its coefficients can be non-trivial, especially for artificial functions such as ReLU.
Figure 7: Fourier spectra of the outputs of (b) ReLU, (c) sigmoid and (d) TDAF functions activated by a sinusoidal wave signal at a dimensionless frequency of 1. Panel (a) shows the spectrum of the reference output produced by the linear identity function. The spectrum of the basic ternariser function is shown in Panel (e).
Consequently, a previous study [72] suggested that two nonlinear physical phenomena can be compared by examining the Fourier spectra of a linear signal processed by the nonlinear systems of interest. Following this approach, we analyse the responses of the ReLU, sigmoid and TDAF activation functions to a purely sinusoidal wave signal at a dimensionless frequency of 1. For the sake of comparison, we also apply the sinusoidal signal to a linear identity function, producing a Fourier spectrum with a sole peak at 1.
The choice of a simple sinusoidal input for analysing the activation function spectra is significant from both machine learning and neuromorphic computing perspectives. It has been shown that a conventional reservoir computing algorithm, which can in principle employ any standard neural activation function [45], is mathematically equivalent to a model that combines the linear weights of the input dataset with its higher-order nonlinear functionals (quadratic, cubic, etc.) [73]. The paper [73] further demonstrates that computational efficiency can be enhanced by designing an algorithm, or using a suitably nonlinear theoretical or physical dynamical system, that sufficiently expands the input over a rich set of nonlinear functionals. This principle was confirmed experimentally using a fluid-mechanical dynamical system as a neuromorphic computer [74]. Within this context, TDAF outperforms both ReLU and sigmoid activations because its Fourier spectrum exhibits all relevant nonlinear harmonics with comparatively higher amplitudes, thereby enabling more effective generation of nonlinear functionals of the input data and resulting in superior computational efficiency [71].
As shown in Fig. 7, the response of ReLU, sigmoid and TDAF to the sinusoidal waves results in the nonlinear generation of higher-order harmonics. According to [72], the strength of the nonlinearity can be quantified using the magnitude of the harmonic peaks and the total number of harmonics produced by the nonlinear process. Focusing on the latter criterion, we can see that ReLU produces the peak at the second-harmonic frequency 2 and then at the fourth and the sixth harmonics. In turn, sigmoid produces the third and fifth peaks only. On the contrary, TDAF generates strong peaks at both odd and even harmonic frequencies. Within the theoretical framework used in this analysis [72], this result suggests that the TDAF exhibits the strongest nonlinearity among the three analysed functions. We also observe that both TDAF and ReLU exhibit DC components in their respective spectra, which further supports our earlier assumption of their similar responses.
Interestingly, the similarity between the Fourier spectra of the sigmoid and ternariser activation functions arises because both introduce strong nonlinearity while preserving a smooth, bounded mapping of input amplitudes. The sigmoid smoothly compresses large inputs toward fixed asymptotic values, effectively saturating the signal and generating higher-order harmonics in its Fourier spectrum. A ternarised function, though discrete, approximates this behaviour by mapping small input variations near zero to a central state (0) and larger magnitudes to $\pm 1$, which retains the overall envelope and shape of the input signal. Consequently, both functions suppress extremely large deviations and concentrate the signal energy around the main frequency components of the original input, leading to a similar harmonic content and spectral distribution. In any case, within this analysis, the ternarisation function exhibits lower nonlinear transformation strength than TDAF.
V-B Probabilistic Ternarisation via Quantum Tunnelling
Using an idealised physical model, we further suggest that the quantum-mechanical effect of QT, underlying tunnel-diode and TDAF operation, naturally enable these electronic devices to function as hardware quantisation units. The electric current through a tunnel diode can be expressed as an energy integral of the quantum transmission probability $T(E,V)$ weighted by the difference of the occupation functions on the two sides of the barrier (see [75, 49, 76] and references therein). In general, we can write
$I(V)\;\propto\;\int_{-\infty}^{\infty}T(E,V)\,\big[f_{p}(E-\mu_{p})-f_{n}(E-\mu_{n})\big]\,\mathrm{d}E\,$, (16)
where $f_{p},f_{n}$ are the Fermi–Dirac distributions on the $p$ and $n$ sides and $\mu_{p}-\mu_{n}=eV$ is the electrochemical potential difference set by the applied bias $V$.
At low temperature, this reduces to an integral over the finite energy window where occupied states on one side overlap empty states on the other side
$I(V)\;\propto\;\int_{\mu_{n}}^{\mu_{p}}T(E,V)\,\mathrm{d}E\,.$ (17)
Hence the voltage dependence of $I(V)$ is determined by how $T(E,V)$ varies with energy and how the bias window $\mu_{n},\mu_{p}]$ sweeps across regions of high/low transmission.
For a simple rectangular barrier (height $V_{0}$, width $a$), the consideration of which suffices for application in the field of machine learning [28], the transmission in the tunnelling regime ($E<V_{0}$) admits the approximate form
$T(E)\;\approx\;\exp\!\big(-2\kappa a\big),\qquad\kappa=\frac{\sqrt{2m(V_{0}-E)}}{\hbar},$ (18)
showing an exponential sensitivity to both $a$ and the barrier offset $V_{0}-E$. Substituting (18) into (17) gives the qualitative behaviour
$I(V)\;\propto\;\int_{\mu_{n}}^{\mu_{p}}\exp\!\big(-2a\sqrt{2m(V_{0}-E)}/\hbar\big)\,\mathrm{d}E\,.$
Two fundamental consequences follow from this relationship. First, regarding the peak current, as $V$ increases from zero, the bias window initially includes energies where $T(E)$ is large (resonant or high transmission), so the integral in (17) and thus $I(V)$ increases. Concerning negative differential resistance, at larger $V$ the overlap of occupied and empty states moves out of the high-$T$ region and into energies where $T(E)$ is smaller; consequently the integral (and $I$) decreases with increasing $V$, producing the $N$-shaped IVI-V curve characteristic of tunnel diodes. If resonant quasi-bound states exist inside the barrier [76, 77], their contribution may be modelled as a narrow resonance $S(E)$ so that the same integral formalism (16) shows how resonances produce sharp peaks in $I(V)$.
Now we return to ternary machine learning models. A common alternative strategy for converting a continuous pre-activation $v$ into a ternary value $\{-1,0,+1\}$ is to use a probabilistic quantisation rule [59]. In this approach, the ternary output is treated as a discrete random variable $W$ with probabilities
$P_{-1}(v)$ $=\Pr(W=-1\,|\,v),$ (19)
$P_{0}(v)$ $=\Pr(W=0\,|\,v),$
$P_{+1}(v)$ $=\Pr(W=+1\,|\,v)\,.$
satisfying $P_{-1}(v)+P_{0}(v)+P_{+1}(v)=1$. One particular method decomposes the ternary choice into two nested Bernoulli decisions [78, 79, 64, 80]. First, we need to decide whether the output is zero or non-zero
$Z\sim\mathrm{Bernoulli}\!\left(p_{0}(v)\right),\qquad\Pr(Z=1\,|\,v)=p_{0}(v)=P_{0}(v)\,.$ (20)
If $Z=1$, the output is set to $W=0$. If $Z=0$, the sign is sampled via a second Bernoulli variable as
$S\sim\mathrm{Bernoulli}\!\left(p_{+}(v)\right),\qquad p_{+}(v)=\frac{P_{+1}(v)}{P_{-1}(v)+P_{+1}(v)}\,.$ (21)
The final ternary output is then
$W=\begin{cases}-1,&Z=0,\,S=0,\\ +1,&Z=0,\,S=1,\\ \;\,0,&Z=1\,.\end{cases}$ (22)
A simpler stochastic quantiser may use a noisy threshold around $v=0$ [67, 59]. Let $\eta$ be a zero-mean random variable (e.g. Gaussian or uniform). Then
$W=\begin{cases}-1,&v+\eta<-\Delta,\\ \;\,0,&|v+\eta|\leq\Delta,\\ +1,&v+\eta>+\Delta,\end{cases}\,$ (23)
which implicitly defines probabilities $P_{-1}(v)=\Pr(v+\eta<-\Delta)$, $P_{0}(v)=\Pr(|v+\eta|\leq\Delta)$ and $P_{+1}(v)=\Pr(v+\eta>\Delta)$.
To employ TDAF in this context, we treat the neural pre-activation $v$ as controlling an effective particle energy $E(v)$ incident on a potential barrier [28]. For each trial, the particle (signal) may be reflected (ternary value 1-1), become trapped or be in temporary bound state [77] (ternary value 0) and transmit through the barrier (ternary value +1+1).
We can model these behaviours with three probabilities $P_{-1}(v)$, $P_{0}(v)$, $P_{+1}(v)$ satisfying
$P_{-1}(v)+P_{0}(v)+P_{+1}(v)=1.$ (24)
Form the physical point of view, for a rectangular barrier of height $V_{0}$ and width $a$, the quantum-mechanical transmission $T(E)$ is the following. For $E<V_{0}$ (tunnelling):
$T(E)=\left[1+\frac{V_{0}^{2}}{4E(V_{0}-E)}\sinh^{2}(\kappa a)\right]^{-1}\,$, (25)
where $\kappa=\dfrac{\sqrt{2m(V_{0}-E)}}{\hbar}$. For $E>V_{0}$ (above-barrier):
$T(E)=\left[1+\frac{V_{0}^{2}}{4E(E-V_{0})}\sin^{2}(ka)\right]^{-1}\,$, (26)
where $k=\dfrac{\sqrt{2m(E-V_{0})}}{\hbar}$.
We can set
$P_{+1}(v)=T(E(v)),$ (27)
and model a temporarily bound state (‘trapped’) probability $S(E)$ as a resonance-like behaviour
$S(E)=s_{0}\exp\!\left(-\frac{(E-E_{r})^{2}}{2\sigma^{2}}\right),$ (28)
where $E_{r}$ is a resonant energy and parameter $\sigma$ controls the width of the resonance lineshape. Renormalisation gives
$P_{+1}=T,\qquad P_{0}=S,\qquad P_{-1}=1-T-S.$ (29)
To interface the dimensionless neural pre-activation $v$ with the quantum-mechanical tunnelling model, we can also map $v$ to an effective carrier energy using the transformation
$E(v)=\alpha v+\beta\,.$ (30)
On physical grounds and based on machine learning parameter fine-tuning, the dimensionless parameters used in the QT ternarisation function are the following. Parameter $\alpha=1.0$ (this is the scaling factor that controls how strongly the neural pre-activation $v$ influences the effective energy $E(v)$; larger $\alpha$ produce sharper transitions between ternary states). The offset $\beta=1.5$ (it sets the reference energy level, determining the alignment of $E(v)$ with respect to the tunnelling barrier). The barrier height $V_{0}=1.51$ and width $a=1.0$ define the rectangular potential barrier, governing the tunnelling probability according to quantum mechanics. Parameters $m=1.0$ and $\hbar=1.0$ are the effective particle mass and reduced Planck constant. Finally, $s_{0}=0.1$, $E_{r}=1.5$ and $\sigma=0.2$ characterise the resonance-like trapped state: $s_{0}$ sets the peak probability of being temporarily captured, $E_{r}$ is the resonant quasi-bound energy and $\sigma$ controls the width of the resonant peak.
We visualise and compare different representations of neural network weights for selected neurons from the first hidden layer of the network, revealing the effect of ternarisation on the weights. Specifically, full-precision weight vectors $W_{1}$ are reshaped into two-dimensional matrices and visually displayed alongside their conventionally balanced ternarised counterparts $btW_{1}$ and QT-ternarised versions $qtW_{1}$. For each selected neuron, the three representations are plotted side by side using greyscale colour maps, allowing a qualitative assessment of the sparsity and quantisation patterns induced by the ternarisation procedures.
In parallel, correlation coefficients between the full-precision and conventionally ternarised weights (approximately 0.85 on average across the neuron population), as well as between full-precision and QT ternarised weights (approximately 0.8), are computed, providing a quantitative measure of how closely the ternarised weights preserve the structure of the full-precision weights.
Obtaining correlation coefficients of this level suggests a moderate degree of linear association, with positive values in one weight distribution generally corresponding to positive values in the other. Thus, both ternarisation approaches moderately reduce the weight resolution and adequately retain structural features of the full-precision weights. In fact, as shown in Table I, a QT-ternarised neural network, with a deliberately reduced number of hidden layers by a factor of two to improve computational speed, confidently reaches 84% classification accuracy.
Figure 8: Visual comparison of weight representations for selected neurons in the first hidden layer. Full-precision weight vectors $W_{1}$ are reshaped into two-dimensional matrices and displayed alongside their conventionally balanced ternarised ($btW_{1}$) and QT-ternarised ($qtW_{1}$) counterparts. For each neuron, the three representations are shown.
Finally, we note that although the software implementation of the QT-based ternarisation formulas may exhibit reduced computational performance relative to conventional ternarisation, the method is well suited for efficient hardware implementation. In analogy with FPGA-based realisations of standard ternarisation, the QT approach can be realised as a compact and low-cost electronic circuit employing tunnel diodes.
This raises the question of whether tunnel diodes can be integrated with FPGAs. Although these two technologies originate from distinct engineering traditions, their integration is not only feasible but technically advantageous for hybrid analogue–digital computing. At the practical level, tunnel diodes can be incorporated as discrete components on a printed circuit board and interfaced with an FPGA through analogue front-end circuitry comprising biasing networks, ADCs and DACs. In this configuration, the tunnel diode provides the high-speed nonlinear transformation required for the operation of the neural network, whereas the FPGA supplies reconfigurable digital control, state updating and post-processing. More sophisticated architectures could embed the tunnel diode within mixed-signal subsystems, allowing the FPGA to dynamically modulate its operating point and gain access to its nonlinear response with high temporal precision. While monolithic co-integration of tunnel diodes with CMOS, such as that used in commercial FPGA fabrication processes, remains limited to proof-of-concept demonstrations due to material and process incompatibilities, research prototypes based on resonant tunnelling diodes integrated with silicon electronics indicate that such hybrid platforms are technologically plausible [81, 82].
VI General discussion
@@
The successful adaptation of the characteristic IVI-V curve of a tunnel diode, and the quantum-tunnelling mechanism that underpins it, as an activation function for neural networks raises interesting questions about the underlying reasons for its improved performance and its potential future applications. TDAF exhibited superior training dynamics, stability and accuracy on the MNIST dataset compared with the sigmoid and ReLU activation functions. Alongside demonstrating the ability of TDAF to function as a quantiser for ternary neural network models, this suggests that TDAF produces a more efficient computing landscape compared with traditional machine learning activation functions.
@@
In practice, a TDAF-based neural network can be implemented using readily available commercial tunnelling [21] and resonant tunnelling [22] diodes. It is noteworthy that systems based on tunnelling diodes consume significantly less power than traditional integrated electronic circuits [33], presenting an opportunity for developing semiconductor chips optimised for power-constrained applications, such as autonomous vehicles, where onboard energy resources are highly limited [10]. Furthermore, neuromorphic computing schemes that exploit negative differential resistance, a key characteristic of tunnelling diodes, have already been proposed [31, 32], suggesting that TDAF functionality could be integrated into existing systems.
Alternatively, a TDAF can be implemented using scanning tunnelling microscopy instrumentation [35]. Amplifiers used in this kind of microscopy operate at microwave frequencies, providing high signal-to-noise ratios and enabling differential resistance spectroscopy measurements [83], which can be used to compute the derivative of TDAF. In low-frequency designs, this derivative can also be obtained using simple analog electronic circuits [84]. Furthermore, QT of individual electrons has been experimentally observed in quantum dots [31], which serve as fundamental building blocks for quantum NC systems and can therefore be integrated with them [11].
Artificial intelligence will be pivotal in future space exploration, but significant technological challenges must be addressed, particularly in energy consumption and thermal management [85]. Modern AI systems rely on efficient cooling to dissipate heat from graphics processing units, yet the vacuum of space renders conventional air-based cooling methods ineffective [85]. This makes heat dissipation a critical concern for spacecraft, satellites and space exploration vehicles, necessitating innovative cooling solutions to ensure the reliable operation of sensitive electronics in extreme environments [86].
@@
VII Conclusion
We introduced a novel activation function, the tunnel-diode activation function (TDAF), derived from the current–voltage (IVI-V) characteristic of a tunnel diode. TDAF demonstrated superior performance compared to two widely used ReLU and sigmoid activation functions, achieving higher accuracy and lower loss during both training and evaluation. Furthermore, we demonstrated that TDAF can be used in ternary neural network models.
We suggested that the improved performance of TDAF can be attributed to the QT effect that underpins the operation of tunnel diodes. This effect introduces greater nonlinearity and expands data into a broader set of nonlinear functionals, enabling machine learning algorithms to process information more efficiently. This physically grounded approach opens new avenues for integrating quantum-inspired mechanisms into machine learning models, bridging the gap between AI and emerging neuromorphic computing platforms.
Future research will focus on further theoretical exploration of the QT-based nonlinearity of TDAF and its implications for network dynamics. Yet, practical implementation of TDAF in hardware-based neuromorphic circuits will be pursued, potentially enabling the development of energy-efficient, high-speed AI systems capable of operating in environments unsuitable for conventional or qubit-based quantum computing hardware.
Comparing the performance of TDAF implementations across different neural network architectures, such as recurrent and convolutional neural networks, reservoir computing systems and transformer-based models, will further test the robustness of the activation function and establish which types of tasks TDAF performs best. Moreover, studying how activation functions affect the loss landscape will provide insight into whether other physically inspired functions may offer advantages for machine learning applications.
As previously mentioned, our results indicate that implementing tunnel diode circuits, which take input data encoded as currents and process it analogously to a feedforward neural network, may yield promising outcomes. Pursuing this research direction could lead to cost-effective AI systems with improved computational efficiency, bridging the gap between quantum-inspired neuromorphic computing systems and practical machine learning implementations.
Data availability
The codes and the datasets used in this research are publicly available and can be accessed at the following links: [57, 62].
Declarations
The authors declare that there is no conflict of interest regarding the publication of this paper.